OpenStudy (anonymous):

−12x2 − 28x + 80 = 0

OpenStudy (anonymous):

-12*2 - 28x +80=0 -24 - 28x + 80= 0 -28x + 56 = 0 56 = 28x x=2

OpenStudy (anonymous):

thank you!

OpenStudy (mathstudent55):

I think @beccal703 misunderstood the problem. −12x2 − 28x + 80 = 0 I think you mean −12x^2 − 28x + 80 = 0 First divide both sides by -4 3x^2 + 7x - 20 = 0 Now factor the left side: (3x - 5)(x + 4) = 0 3x - 5 = 0 or x + 4 = 0 3x = 5 or x = -4 x = 5/3 or x = -4